A computer was sold for 1,350 euro + tax at 20% percent. The rate of tax was 21%percent.Find the total price of the computer.
The price of a camcorder, including tax at 20%percent is 528 euro. Find the price of the camcorder before tax is added.
What sum of money will earn 45 euro simple interest if it is invested at 9% percent per year for 1 year.

Answer by rfer(12670)

1)
1350*1.2=1620
1350*1.21=1633.50
2)
528/1.2=440
3)
45/.09=500